# One Model For All: Partial Diffusion for Unified Try-On and Try-Off in Any Pose

### v1

@@ A.2 Evaluation Metrics
We provide the model image, garment image, and the try-on result as input to GPT-4o-mini. The prompt for the Multimodal Large Language Model, as presented in Fig. 7.
@@
Appendix B More Qualitative Results
B.1 Virtual Try-on
Person-to-person Virtual Try-on.
Fig. 8 presents additional try-on comparative results in the person-to-person scenario on the VITON-HD dataset. Specifically, when the input clothing is not an exhibition garment, the input warp cloth often exhibits incomplete contours and distorted textures, which further exacerbates the artifacts in the try-on results, ultimately leading to suboptimal performance. Fig. 9 and Fig. 10 present more generation results of our proposed OMFA model in this task, further demonstrating that our model maintains excellent detail preservation in both the try-off and try-on steps, leading to robust and high-fidelity results.
@@ B.2 Vitual Try-off
Appendix C Limitations
In our experiments, we observed that our method fails to preserve regions beyond the person’s identity and the target garment, such as the background and lower-body cloth. This issue mainly stems from the lack of constraints on these regions in our mask-free approach. How to preserve such non-garment areas without relying on explicit masks is left as a direction for future work.

### v3

@@ 7.2 Evaluation Metrics
We provide the model image, garment image, and the try-on result as input to GPT-4o-mini. The prompt for the multimodal large language model is presented in Fig. 9.
8 Additional Ablation Studies
Experimental Setup for SMPL-X Ablation.
Specifically, we trained two low-resolution models on VITON-HD, using DensePose (provided by the VITON-HD preprocessing) and SMPL-X respectively as structural conditioning, which were concatenated along the input channels. The configuration uses a resolution of 384×512, a per-GPU batch size of 8, and two 80-GB A100 GPUs for 45,000 training steps. At inference, the guidance scale is set to 2.0.
Ablation Studies for Joint Training Strategy.
To evaluate the impact of unified training on performance, we conduct an ablation using the same low-resolution configuration as described above. In the joint-training setting, each batch is duplicated and divided into two halves, one for try-off training and the other for try-on training. Tab.  6 compares two training setups with no performance degradation. We attribute this to our bidirectional Tweedie diffusion mechanism, which explicitly defines the context and generation task by adjusting the combination of noising targets and conditions, thereby avoiding cross-task context confusion. Moreover, unified training encourages the model to learn bidirectional, shared feature correlations between garments and the human body, yielding more essential representations and improving the performance and robustness of both tasks.
Table 6: Quantity ablation of unified training strategy. where “(u)” indicates that the metric is computed in the unpaired setting. Training the network to learn both try-on and try-off tasks does not degrade performance.
[…]
Try-on training 0.840 0.150 0.866 11.919 3.603
Joint training 0.839 0.152 0.872 11.781 3.602
9 More Qualitative Results
9.1 Virtual Try-on
Person-to-person Virtual Try-on.
Fig. 10 presents additional try-on comparative results in the person-to-person scenario on the VITON-HD dataset. Specifically, when the input clothing is not an exhibition garment, the input warp cloth often exhibits incomplete contours and distorted textures, which further exacerbates the artifacts in the try-on results, ultimately leading to suboptimal performance. Fig. 11 and Fig. 12 present more generated results of our proposed OMFA model in this task, further demonstrating that our model maintains excellent detail preservation in both the try-off and try-on steps, leading to robust and high-fidelity results.
Multi-pose Virtual Try-on.
9.2 Virtual Try-off
9.3 User Study
We conducted a user study with 50 participants using the model trained on the VITON-HD dataset. Each participant was randomly assigned 10 samples from a pool of 50 for evaluation, with each sample containing six different virtual try-on results generated in the person-to-person scenario. Participants were asked to choose the best result using three criteria: image fidelity, human identity, and garment consistency. We totaled the number of times each method was chosen as the best across all test samples and calculated the average voting proportion for each method. As shown in Tab. 7, our method had the highest average voting proportion among all examples, indicating visually superior results and a significant advantage in human evaluation.
Table 7: User study results. We report the best-choice rate for our method and seven other methods, including StableVITON, OOTDiffusion, and CatVTON.
OMFA IDM-VTON MV-VTON others
Best Choice Rate 46% 8.6% 18.8% 26.6%
10 Limitations
Due to a lack of paired data for multi-layer garments, our proposed method does not provide multi-layer try-on/try-off. Furthermore, our architecture is only intended for a single garment input, whereas multiple garment inputs may dramatically extend the input sequence. In the future, we will incorporate more in-the-wild data to develop computationally efficient virtual try-on solutions that are more in line with real-world application scenarios.